OpenStudy (anonymous):

solve using substitution or the linear combination/elimination method 5x-25y=40 x=4y+6 thx

OpenStudy (radar):

x=4y+6 5(4y+6)-25y=40 20y+30-25y=40 -5y=10 y=-2

OpenStudy (radar):

that solved for y now substitute the y value in either equation and solve for x. Then check answer.

OpenStudy (anonymous):

thanks what method did you use?

OpenStudy (radar):

I used the substitution method.

OpenStudy (radar):

In the 2nd line you will note that the x has been substituted with 4y+6/

OpenStudy (radar):

Looking at the 2nd equation,x=4y+6 is a good one to use to solve for x x=4(-2)+6=-2..........my goodness they are both -2. Better check and verify. -2= -8+6 -2=-2 thats good, now for the first one 5(-2) -25(-2)=40 -10+50=40 40=40 good, it is correct.

OpenStudy (radar):

be back in a few

OpenStudy (anonymous):

thank you for checking sir, I came up with that answer also and I appreciate you double checking.

OpenStudy (anonymous):

ok thank you could you please help me with this one:7x+2y=10 3x-5y=-25

OpenStudy (radar):

I'm back, you still here.

OpenStudy (anonymous):

yes

OpenStudy (radar):

Is there a preference in the method to use to solve the new problem?

OpenStudy (anonymous):

either substitution or the linear combination/elimination method thank you for asking

OpenStudy (radar):

I will use the elimination method, to you this method, you want to get one of the variables ( x or y) to have equal and opposite coefficients. I choose to eliminate the y to do that I must multiply all terms in the first equation by this will give me: 21x+6y=30. Now I multiply 2nd equation by a -7 (all terms) getting: -21x +35y=175 I now add the two equations and eliminate the x -------------- 41y=205 y=5 we now solve for x by subsstituting the value of y in your equation, either one.

OpenStudy (radar):

*eliminate the "x" made a typo

OpenStudy (anonymous):

would x =0 ?

OpenStudy (radar):

Lets see.. I will use 7x+2y=10 7x + 2(5)=10 7x=0 x=0 Yes you're correct.

OpenStudy (anonymous):

thank you, I am trying to study for a semester exam, would you be willing to help me with a few more? I really appreciate your help

OpenStudy (radar):

I will have to run, Thurs night is my TV night lol, Missing, Person of Interest, and The Mentalist.

OpenStudy (radar):

Post again, there is usually someone to help.

OpenStudy (anonymous):

ok thank you

OpenStudy (radar):

Good luck with the semester exam, I feel that you are going to do OK on that exam.

OpenStudy (anonymous):

alright i appreciate your help